# Breaking degeneracies in exoplanetary parameters through self-consistent atmosphere-interior modelling

### v2

@@ 2.4 Model comparison
An overview of features available in models currently used can be found in table 3. Marley et al. (2021) is the first version of a series of papers, Karalidi et al. (2021) is the second paper in this series and adds disequilibrium chemistry. Clouds are planned in a latter version. The models are for the non-irradiated case, they are developed for emission spectra and do not provide transit spectra. The current available grid by Marley et al. (2021) uses the SCVH (Saumon et al. 1995) parametric equation of state instead of the more recent ab-initio equations of state used in Phillips et al. (2020).
Using the public evolution grid from Phillips et al. (2020) as well as from Marley et al. (2021), we can compare the results given by our models for the non-irradiated case. Figure 2 illustrates a comparative analysis of planetary models within the 1-5 MJ range. We conducted this comparison using both the ab-initio EOS (Chabrier et al. 2019) and the SCVH EOS (Saumon et al. 1995), aiming to provide a comprehensive insight into the variations between these models.
It is noteworthy that when directly comparing the calculations (non-interpolated), the disparities between the models and our findings remain modest, staying below in most of the evaluated range. Interestingly, our results consistently exhibit slightly larger radii at 200 to 400K, regardless of the EOS used. Above 400K, our outcomes align closely with those of Phillips et al. (2020). However, when considering lower masses, a noticeable divergence emerges with Marley et al. (2021) as temperatures increase. While differences in atmosphere models could contribute to the observed variations, it is important to acknowledge their relatively minor impact compared to the substantial discrepancies arising from distinct EOS selections. For the remainder of this work, the ab-initio EOS for effective temperatures above 400K is used.
@@
In our verification process, we assess the reliability of the interpolation tool by comparing our interpolated results with those of Phillips et al. (2020), utilizing a dataset computed at masses distinct from those within our target grid. Remarkably, the interpolation yields highly comparable results and the righter most graphs of Figure 2 are comparable to the same equivalent graphs on the left using actual grid points.
The previous comparison shows the overall validity of our model compared to currently available models, this for the two different equations of state tested here. While this comparison is done in the non-irradiated case, there is no difference in the methodology used when adding stellar irradiation. The capacity to interpolate within the grid allows for a more flexible usage and reduced overall computation time when evaluating planetary parameters as is done throughout this present work.
@@ 3 Stellar irradiation
Considering stellar irradiation, we can define the effective temperature of a planet as
With Tirr the equivalent irradiation temperature and Abond the Bond Albedo.
Planets cool over time by evacuating internal energy into space and reducing their entropy (S). This is summarized in equation 7.
@@
For the rest of this work, the Bond albedo () derived from Exo-REM
@@
is approximately , assuming a M star for the light source (BT-Settle spectrum at 3500K from Allard et al. (2012)) and a cloud-free atmosphere. As such, for irradiated planets, the equilibrium temperature is approximately equivalent to the irradiation temperature.
@@
In the context of a given pressure-temperature profile for a planet, distinct radiative and convective regions become discernible. The location of the radiative-convective boundary, which demarcates these regions, hinges upon the equilibrium between stellar irradiation and the internal flux associated with planetary cooling. The result of this is shown in Figure 1, where an array of profiles, each characterized by varying irradiation temperatures, is depicted. Heightened irradiation extends the reach of the radiative domain to higher pressures within the atmosphere, a trend consistently revealed in Figure 3 within the TT domain. This same figure shows how the radiative-convective equilibrium pressure (RCEP) evolves as a planet cools. For non-irradiated planets, there is little to no evolution of the RCEP with intrinsic temperature. Above approximately 400K in Tirr, there is a significant burying of the RCEP for decreasing intrinsic temperatures, equivalent to a planet ageing. This result helps explain why non-irradiated models are valid for fixed linkage pressures between atmosphere and interior and why irradiated models need variability in this pressure. As the RCEP is driven to lower strata due to irradiation, the validity of certain hypothesis of atmosphere models is put into question. Notably when the atmosphere model relies on the ideal gas law, significant deviation can arise at pressures above 103 bars. While we do not treat this in this current work it remains important to insure that the density does not deviate too greatly between interior and atmosphere model. This can be the case for very old irradiated planets with T150K which Figure 4 shows to be a finely populated region considering current observations. Furthermore, models with fixed gravity throughout the atmosphere, can suffer from a deviations of up to 10 in gravity between the top and bottom of the atmosphere (1bar to RCEP). The adaptive linking implimented in this work accounts for this variability as was discussed in section 2.3.
@@
As the heat penetrates further into the atmospheric layers, it notably elevates the entropy of the selected adiabatic gradient for the internal structure. Consequently, this upshifted gradient engenders a seemingly decelerated cooling rate for the planet. Thus emerges the need to establish an intricate relationship between age, mass, and radius within an irradiated evolution model for Jupiter-like exoplanets.
@@
The self-coherent linking of the gravity between the atmosphere model and the interior model, means that equation 7 needs to be solved on a non regular grid in entropy. To complete this, equation 7 can be rewritten as equation 8 by considering that the contribution from the isothermal region of the core, where the entropy is non constant, is negligible. This is shown in Mordasini et al. (2012a), the cores luminosity is negligible during the evolution phase of a planet when compared to the internal luminosity. Mordasini et al. (2012a) defines the internal luminosity as :
@@
We solve equation 8 using a Runge Kutta 4 numerical method along an interpolated regular grid in Tint for a given set of planetary parameters. We use the specific entropies of young planets under hot start assumption from Marley et al. (2007) as initial entropies. A rigorous exploration of the impact of the initial entropy should be undertaken for the case of young giant planets whose ages are estimated to be below the Kelvin-Helmholtz timescale. Using equation 11 from Guillot & Gautier (2015), we find a timescale of for 51 Eridani b, using the best retrieval from section 7.
@@
In their study, Thorngren et al. (2019) established a connection between stellar irradiation and intrinsic temperature, to explain the radius inflation of hot Jupiters, by building upon their earlier research, Thorngren & Fortney (2018). This earlier work had already established a link between the heating efficiency of incoming stellar radiation and its role in understanding the observed radius of a group of transiting planets. They use the Thorngren et al. (2016) mass-metallicity relationship to further increase the Bayesian heating efficiency fit. They propose a Gaussian fit between the model based intrinsic temperature and the measured equilibrium temperature and radius. We chose to derive, via interpolation, intrinsic temperatures values for planets within the NASA Exoplanet archive, taking into account the mass-metallicity trend identified in Thorngren et al. (2016).
@@
Figure 4 shows how the intrinsic temperature varies with observables. The intrinsic temperature values for each planet are obtained by interpolating from the grid of models. We identify a link between the intrinsic temperature and the equilibrium temperature, Figure 4.a shows that increased irradiation leads to an increased intrinsic temperature following a quasi-linear trend given by equation 10.
@@
This can from a modelling perspective be seen as a warming of the internal adiabat. Higher mass planets have a wider spread of intrinsic temperatures, this is to be expected as larger planets cool at a slower rhythm. Figure 4.b shows that the mass of the planet plays a role. Higher mass planets seem to show a larger spread in intrinsic temperatures compared to lower mass planets. This can by understood as higher mass planets taking longer to evacuate intrinsic energy. Figure 4.c shows that the intrinsic temperature that we derive is linked to the stellar age. We would expect that a planet with a lower intrinsic temperature would be older. We observe that planets with higher irradiation temperatures maintain higher intrinsic temperatures. Finally Figure 4.d shows that the intrinsic temperature increases for more inflated planets. There is a potential plateau that appears around the 400K region, where the radius is more sensitive to the intrinsic temperature. The stand out samples far above the 400K intrinsic temperature region should be further studied.
@@
Compared to Thorngren et al. (2019) we observe no significant change in trend in the intrinsic flux for an increasing irradiation. In addition, the derived values for the intrinsic temperature are generally lower than in Thorngren et al. (2019). On the lower end of the irradiation temperature, there is relatively good agreement between both results. However, as is shown by Figure 4.a, the majority of planets have a predicted intrinsic temperature between 200 and 400K which is as low as half of that which is shown in the fit by Thorngren et al. (2019). This result corresponds to a heat transfer of to the interior for a HD209458b-like planet (Henry et al. 2000) which is in agreement with results from Guillot & Showman (2002) for the flux transfer required to explain the radius of HD209458b taking into account energy dissipation in the interior. Compared to the efficiency of heat transfer required in Thorngren et al. (2019), this makes the observed inflated radii easier to reproduce by physical processes such as ohmic dissipation (Batygin & Stevenson 2010) or advection of heat by the atmospheric circulation (Tremblin et al. 2017). We believe that equation 10 should serve as an approximate temperature to be used when running global climate models for hot Jupiter’s as is done for instance in Komacek et al. (2022).
@@ 4 Breaking degeneracies
To assess planetary parameters based on observable data, a commonly employed method is a statistical approach, such as Bayesian inference. Nevertheless, when faced with limited observables, it becomes intriguing to explore whether it is feasible to resolve parameter degeneracies using physical models. In this context, we will focus on two specific parameters: metallicity and core size. Our objective is to investigate whether discrepancies in these parameters, within the (Tint; Tirr) space at a fixed mass, result in discernible variations in planetary radii.
@@ 4.1 Metallicity
By linking the molecular mass of the atmosphere in the convective region with the helium fraction of the interior, we can adjust the interior structure to take into account varying metallicities. We can hence compare the TT space between different metallicities and uncover regions where the irradiation, age and radius are sufficient variables to distinguish the planets metallicity.
Figure 5.a illustrates that a planet with 10 times the solar metallicity, with a very high intrinsic temperatures, may exhibit a larger radius than an equivalent planet with 1 times solar metallicity. In contrast, the right panel shows that a planet with 30 times the solar metallicity consistently maintains a smaller radius across the range of intrinsic and irradiation temperatures. The increase in the radius, and subsequent decrease with metallicity, is better shown in the lower panel of Figure 5. As we increase the temperature, the radius inflates for all cases as expected. When increasing metallicity at fixed temperature, the radius also increases, but beyond a certain metallicity (0.5) the radius decreases again. We hence observe two regions, one where an increasing metallicity leads to an increase in absorption and hence the thermal effects dominate. A second region where the molecular mass dominates and gravity forces the radius to contract. We therefore observe a ”C” shape in the (Z;Tint) space. Thus, for a fixed radius, higher metallicities necessitate a lower intrinsic temperature because the radiative-convective boundary occurs at a lower pressure due to the increased opacity and greenhouse effect. However, beyond a certain metallicity threshold, the effect of the mean molecular mass on the atmospheric scale height dominates the effect of the opacity. Hence higher intrinsic temperatures are required to maintain the radius and counteract the effects of gravity. The threshold is temperature dependent, 0.2 dex for 200K and 1 dex for 800K.
@@ 4.2 Core size
Exploring the properties of cores within gaseous planets is a topic of significant interest. Fundamental questions persist regarding their distribution within these planets and their contribution to both mass and radius. The layout of the core and even the presence of one, is a topic of significant interest when considering formation mechanisms (Pollack et al. 1996; Mordasini et al. 2012b, c). However, when studying giant exoplanets, the primary question that arises is whether variations in core characteristics have a discernible impact on observable parameters, such as the planet’s radius. To address this inquiry, we introduce an additional dimension to our analysis by considering different core sizes, from 1 to 40 times Earth mass (ME).
@@
Interestingly, we observe a weaker effect on the radius with increasing irradiation temperature compared to increasing intrinsic temperature. Consequently, the core of young planets has the potential to become a quantifiable parameter, so long as they have a sufficiently high intrinsic temperature. Exploring this realm of the parameter space is poised to offer valuable insights into alternative interior designs, such as those featuring diluted cores.
@@ 5 Grid interpolation and retrieval
As discussed in the model verification and shown in Figure 2.c, it is possible to obtain the modeled radius of a planet at any point within the grid. Hence the radius is a continuous parameter. In order to efficiently develop grid retrieval tools, it is equally interesting to be able to interpolate the spectra (transit or emission) of a simulated planet with a given set of input parameters, defined as follows :
It is not possible to do this by simply using the LinearNDInterpolator interpolator as for the radius. The issue is that this will lead to leakage between different wavelengths and the resulting spectrum will be heavily influenced by the distance between the new wavelength as well as the other planetary parameters. As such a solution is to create an array of interpolators, following :
For each wavelength of the spectra generated (), we create a unique interpolation function which has for input the planetary parameters. It is no longer possible to leak information from one wavelength into another as the unique interpolators are independent from one another. The draw back of this method is that it is essential to reinterpolate each point of grid to a desired wavelength range before obtaining a new spectrum for a given planet.
Utilizing Markov Chain Monte Carlo (MCMC) sampling on a continuous grid with the emcee
python library, we aim to ascertain the optimal planetary parameters that can elucidate an observed spectrum. To evaluate this approach, we introduce Gaussian noise () to a spectrum extracted from the grid, hence not interpolated.
The selected spectrum pertains to an artificial planet for which we fixed the parameters to 1 Jupiter mass, featuring intrinsic and irradiation temperatures of 800K and 100K respectively, alongside a solar metallicity and a 10 ME core. The resolution used is chosen so as to prepare for the test case in section 7 (). By considering a H-band spectrum and irradiation temperature as observable factors, our assessment involves sampling spectrum while treating mass, intrinsic temperature, metallicity and core mass as variables. We consider that the planet is placed at 10 parsecs from the observer. We adjust the flux level to account for this and the model radius using :
@@
We consider two distinct terms, the first consisting of the double sum. This takes into account the fact that a spectrum can be a mosaic of spectra from multiple instruments. In this case, we propose allowing each instrument to be treated as an equally weighted observable. The second term is the simple sum which takes into account one or more physical characteristics of the planet. For instance a mass measurement using radial velocity or a radius using the average transit depth. The proposed term is given to allow for a greater importance to be given to certain observed characteristics.
For the example studied here, equation 14 is reduced to the first term. We use uniform priors except at the grid limits where we impose an infinite barrier. The priors are chosen as follows :
The outcome of this evaluation is illustrated in annex 12, which presents a corner plot of the MCMC-based interpolated grid retrieval. The radius distribution is added by interpolating the four other parameters (Mp,Tint,Z,Mc) for each walker and iteration. Figure 7 shows the retrieved interpolated spectrum as a comparison to the degraded and original spectra.
The results indicate a good capacity to retrieve a spectrum using interpolators and MCMC sampling. They reveal a relatively accurate retrieval of the intrinsic temperature, metallicity and mass. The core mass is retrieved correctly however with a relatively broad posterior distribution. This indicates that future core mass estimations will at best be given with around a 5 ME uncertainty for a spectrum with similar characteristics as the one shown in Figure 7.

### v3

@@ 2.4 Model comparison
An overview of the features available in the models used here can be found in Table 3. Marley et al. (2021) is the first of a series of papers, and Karalidi et al. (2021) is the second paper in this series and adds disequilibrium chemistry. Clouds can be found in the latter Morley et al. (2024) model. The models are for the non-irradiated case; they are developed for emission spectra and do not provide transit spectra.
Using the public evolution grid from Phillips et al. (2020), as well as that from Marley et al. (2021), we can compare the results given by our models for the non-irradiated case. Figure 2 illustrates a comparative analysis of planetary models within the 1-5 MJ range. We conducted this comparison using both the ab initio EOS (Chabrier et al. 2019) and the SCVH EOS (Saumon et al. 1995), aiming to provide a comprehensive description of the variations between these models.
It is noteworthy that when directly comparing the calculations (non-interpolated), the disparities between the models and our findings remain modest, and below in most of the evaluated range. Interestingly, our results consistently exhibit slightly larger radii at 200 to 400K, regardless of the EOS used. Above 400K, our outcomes align closely with those of Phillips et al. (2020). However, when considering lower masses, a noticeable divergence emerges between our findings and those of Marley et al. (2021) as temperatures increase. While differences in atmosphere models could contribute to the observed variations, it is important to acknowledge their relatively minor impact compared to the substantial discrepancies arising from distinct EOS selections. For the remainder of this work, we use the ab initio EOS for effective temperatures above 400K.
@@
In our verification process, we assess the reliability of the interpolation tool by comparing our interpolated results with those of Phillips et al. (2020), using a dataset computed at masses distinct from those within our target grid. Remarkably, the interpolation yields highly comparable results and the rightmost graphs of Figure 2 are comparable to the equivalent graphs on the left, which use actual grid points.
The above comparison shows the overall validity of our model compared to currently available models, this for the two different EOSs tested here. While this comparison is made for the non-irradiated case, there is no difference in the methodology used when adding stellar irradiation. The capacity to interpolate within the grid allows for more flexible use and reduced overall computation time when evaluating planetary parameters as done throughout the present work.
@@ 3 Stellar irradiation
Considering stellar irradiation, we can define the effective tem- perature of a planet as
Planets cool over time by evacuating internal energy into space and reducing their entropy (). This is summarised by the following equation,
@@
For the remainder of this work, the Bond albedo () derived from Exo-REM
@@
is approximately zero, assuming an M star for the light source (BT-Settle spectrum at 3500K from Allard et al. (2012)) and a cloud-free atmosphere. As such, for irradiated planets, the equilibrium temperature is approximately equivalent to the irradiation temperature.
@@
In the context of a given pressure–temperature profile for a planet, distinct radiative and convective regions become discernible. The location of the radiative-convective boundary, which demarcates these regions, hinges upon the equilibrium between stellar irradiation and the internal flux associated with planetary cooling. The result of this is shown in Figure 1, where an array of profiles is depicted, each characterised by a different irradiation temperature. Heightened irradiation extends the reach of the radiative domain to higher pressures within the atmosphere, a trend consistently revealed in Figure 3 within the TT domain. This same figure shows how the radiative–convective equilibrium pressure (RCEP) evolves as a planet cools. For non-irradiated planets, there is little to no evolution of the RCEP with intrinsic temperature. Above approximately 400K in Tirr, there is significant burying of the RCEP for decreasing intrinsic temperatures, equivalent to a planet ageing. This result helps explain why non-irradiated models are valid for fixed linkage pressures between atmosphere and interior and why irradiated models need variability in this pressure. As the RCEP is driven to lower strata due to irradiation, the validity of certain assumptions regarding atmosphere models is put into question. Notably when the atmosphere model relies on the ideal gas law, significant deviation can arise at pressures above 103 bars. While we do not treat this in the present work, it remains important to ensure that the density does not deviate too greatly between the interior model and the atmosphere model. This can be the case for very old irradiated planets with T150K, which Figure 4 shows to be a sparsely populated region considering current observations. Furthermore, models with fixed gravity throughout the atmosphere can suffer from deviations of up to 10 in gravity between the top and bottom of the atmosphere (1bar to RCEP). The adaptive linking implemented in this work accounts for this variability, as discussed in Section 2.3.
@@
As the heat penetrates further into the atmospheric layers, it notably elevates the entropy of the selected adiabatic gradient for the internal structure. Consequently, this upshifted gradient engenders a seemingly decelerated cooling rate for the planet. This is where a requirement emerges to establish an intricate relationship between age, mass, and radius within an irradiated evolution model for Jupiter-like exoplanets.
@@
The self-coherent linking of the gravity between the atmosphere model and the interior model requires that we solve equation 7 on an irregular grid in entropy. To complete this, equation 7 can be rewritten as equation 8 by considering a negligible contribution from the isothermal region of the core, where the entropy is variable. This is shown in Mordasini et al. (2012); the luminosity of the core is negligible during the evolution phase of a planet when compared to the internal luminosity. Mordasini et al. (2012) defines the internal luminosity as
@@
We solve equation 8 using a Runge Kutta 4 numerical method along an interpolated regular grid in Tint for a given set of planetary parameters. We use the specific entropies of young planets under the assumption of a hot start from Marley et al. (2007) as initial entropies. A rigorous exploration of the impact of the initial entropy should be undertaken for the case of young giant planets whose ages are estimated to be below the Kelvin-Helmholtz timescale. Using equation 11 from Guillot & Gautier (2015), we find a timescale of for 51 Eridani b using the best retrieval from section 7.
@@
In their study, Thorngren et al. (2019) established a connection between stellar irradiation and intrinsic temperature to explain the radius inflation of hot Jupiters by building upon their earlier research (Thorngren & Fortney 2018). This earlier work established a link between the heating efficiency of incoming stellar radiation and its role in our understanding of the observed radius of a group of transiting planets. These authors used the Thorngren et al. (2016) mass–metallicity relationship to further increase the Bayesian heating efficiency fit. They propose a Gaussian fit between the model-based intrinsic temperature and the measured equilibrium temperature and radius. We chose to derive, via interpolation, intrinsic temperature values for planets within the NASA Exoplanet archive, taking into account the mass–metallicity trend identified in Thorngren et al. (2016).
@@
Figure 4 shows how the intrinsic temperature varies with observables. The intrinsic temperature values for each planet are obtained by interpolating from the grid of models. We identify a link between the intrinsic temperature and the equilibrium temperature. Figure 4(a) shows that increased irradiation leads to an increased intrinsic temperature following a quasi-linear trend given by equation 10.
@@
From a modeling perspective, this can be interpreted as a warming of the internal adiabat. Higher-mass planets exhibit a broader range of intrinsic temperatures, as shown in Figure 4(b), likely due to their slower cooling rates and longer timescales for evacuating intrinsic energy compared to lower-mass planets. Figure 4(c) shows that the intrinsic temperature that we derive is linked to stellar age. We would expect that a planet with a lower intrinsic temperature would be older. We observe that planets with higher irradiation temperatures maintain higher intrinsic temperatures. Finally, Figure 4(d) shows that the intrinsic temperature increases for more inflated planets. There is a potential plateau that appears around the 400K region, where the radius is more sensitive to the intrinsic temperature. The stand-out samples far above the 400K intrinsic temperature region should be studied further.
@@
COmparing our results with those of Thorngren et al. (2019), we observe no significant change in the trend of the intrinsic flux with increasing irradiation. In addition, the derived values for the intrinsic temperature are generally lower than those found by these authors. On the lower end of the irradiation temperature range, there is relatively good agreement between our results and theirs. However, as shown in Figure 4(a), the majority of planets have a predicted intrinsic temperature of between 200 and 400K, which is as low as half of that shown in the fit by Thorngren et al. (2019). This result corresponds to a heat transfer of to the interior for a HD209458b-like planet (Henry et al. 2000), which is in agreement with the results of Guillot & Showman (2002) for the flux transfer required to explain the radius of HD209458b, taking into account energy dissipation in the interior. Compared to the efficiency of heat transfer required in Thorngren et al. (2019), this makes the observed inflated radii easier to reproduce by physical processes such as ohmic dissipation (Batygin & Stevenson 2010) or advection of heat by the atmospheric circulation (Tremblin et al. 2017). We believe that equation 10 should serve as an approximate temperature to be used when running global climate models for hot Jupiters, as is done for instance in Komacek et al. (2022).
@@ 4 Breaking degeneracies
To assess planetary parameters based on observable data, a commonly employed method is a statistical approach, such as Bayesian inference. Nevertheless, when faced with limited observables, it becomes intriguing to explore whether it is feasible to resolve parameter degeneracies using physical models. In this context, we focus on two specific parameters: metallicity and core size. Our objective is to investigate whether discrepancies in these parameters, within the (Tint; Tirr) space at a fixed mass, result in discernible variations in planetary radii.
@@ 4.1 Metallicity
By linking the molecular mass of the atmosphere in the convective region with the helium fraction of the interior, we can adjust the interior structure to take into account varying metallicities. We can then compare the TT space between different metallicities and uncover regions where the irradiation, age, and radius are sufficient variables to distinguish the planet’s metallicity.
Figure 5(a) illustrates that a planet with ten times the solar metallicity and a very high intrinsic temperature may exhibit a larger radius than an equivalent planet with solar metallicity. In contrast, the right panel shows that a planet with 30 times the solar metallicity consistently maintains a smaller radius across the range of intrinsic and irradiation temperatures. The increase and subsequent decrease in the radius with metallicity is better shown in the lower panel of Figure 5. As we increase the temperature, the radius inflates for all cases, as expected. When increasing metallicity at fixed temperature, the radius also increases, but beyond a certain metallicity (0.5) the radius decreases again. We therefore observe two regions, one where an increasing metallicity leads to an increase in absorption and hence the thermal effects dominate, and a second region where the molecular mass dominates and gravity forces the radius to contract. We therefore observe a ‘C’ shape in the (Z;Tint) space. Thus, for a fixed radius, higher metallicities necessitate a lower intrinsic temperature because the radiative–convective boundary occurs at a lower pressure due to the increased opacity and greenhouse effect. However, beyond a certain metallicity threshold, the effect of the mean molecular mass on the atmospheric scale height dominates the effect of the opacity. Hence, higher intrinsic temperatures are required to maintain the radius and counteract the effects of gravity. The threshold is temperature dependent, 0.2 dex for 200K and 1 dex for 800K.
@@ 4.2 Core size
Exploring the properties of cores within gaseous planets is a topic of significant interest, and fundamental questions persist regarding their distribution within these planets and their contribution to both mass and radius. The layout of the core and even the presence of one, is a topic that has received significant attention where formation mechanisms are considered (Pollack et al. 1996; Mordasini et al. 2012). However, when studying giant exoplanets, the primary question that arises pertains to whether variations in core characteristics have a discernible impact on observable parameters, such as the planet’s radius. To explore this topic, we introduced an additional dimension to our analysis by considering different core sizes, from 1 to 40 times Earth mass (ME).
@@
Interestingly, we observe a weaker effect on the radius with increasing irradiation temperature compared to increasing intrinsic temperature. Consequently, the core of young planets has the potential to become a quantifiable parameter, so long as they have a sufficiently high intrinsic temperature. Exploration of this realm of the parameter space is poised to offer valuable insights into alternative interior designs, such as those featuring diluted cores.
@@ 5 Grid interpolation and retrieval
As discussed Section 2.4 and shown in Figure 2 (c), it is possible to obtain the modelled radius of a planet at any point within the grid. Hence the radius is a continuous parameter. In order to efficiently develop grid retrieval tools, it is equally interesting to be able to interpolate the spectra (transit or emission) of a simulated planet with a given set of input parameters, defined as follows:
It is not possible to do this simply using the LinearNDInterpolator interpolator as for the radius. The issue is that this will lead to leakage between different wavelengths and the resulting spectrum will be heavily influenced by the distance between the old and new wavelengths, as well as the other planetary parameters. One solution is to create an array of interpolators, following:
For each wavelength of the generated spectra (), we create a unique interpolation function that takes the planetary parameters as input. It is no longer possible to leak information from one wavelength into another, as the unique interpolators are independent of one another. The drawback of this method is that it is essential to re-interpolate each point on the grid to a desired wavelength range before obtaining a new spectrum for a given planet.
Using Markov Chain Monte Carlo (MCMC) sampling on a continuous grid with the emcee
Python library, we aim to ascertain the optimal planetary parameters that can elucidate an observed spectrum. To evaluate this approach, we introduce Gaussian noise () to a spectrum extracted from the grid, and therefore not interpolated.
The selected spectrum pertains to an artificial planet for which we fixed the parameters to 1 Jupiter mass, featuring intrinsic and irradiation temperatures of 800K and 100K, respectively, alongside solar metallicity and a 10 ME core. The resolution used is chosen so as to prepare for the test case in Section 7 (). By considering a H-band spectrum and irradiation temperature as observable factors, our assessment involves sampling the spectrum while treating the mass, intrinsic temperature, metallicity, and core mass as variables. We consider that the planet is placed at 10 parsecs from the observer. We adjust the flux level to account for this and the model radius using:
@@
We consider two distinct terms, the first consisting of the double sum. This takes into account the fact that a spectrum can be a mosaic of spectra from multiple instruments. In this case, we propose to allow each instrument to be treated as an equally weighted observable. The second term is the simple sum that takes into account one or more physical characteristics of the planet, such as a mass measurement using radial velocity or a radius using the average transit depth. The proposed term is given to allow greater importance to be given to certain observed characteristics.
For the example studied here, equation 14 is reduced to the first term. We use uniform priors, except at the grid limits where we impose an infinite barrier. The priors are chosen as follows:
The outcome of this evaluation is illustrated in Appendix 12, which presents a corner plot of the MCMC-based interpolated grid retrieval. The radius distribution is added by interpolating the four other parameters (Mp,Tint,Z,Mc) for each walker and iteration. Figure 7 shows the retrieved interpolated spectrum compared to the degraded and original spectra.
The results indicate a good capacity to retrieve a spectrum using interpolators and MCMC sampling; they reveal a relatively accurate retrieval of the intrinsic temperature, metallicity, and mass. The core mass is retrieved correctly, albeit with a relatively broad posterior distribution. This indicates that future core-mass estimations will at best be given with around a 5 ME uncertainty for a spectrum with similar characteristics to the one shown in Figure 7.